# Layer-Dependent Quantum Anomalous Hall Effect in Rhombohedral Graphene

### v1

@@ Abstract
The quantum anomalous Hall (QAH) effect, first proposed by the Haldane model, has become a paradigmatic example of application of band topology to condensed matter physics. The recent experimental discoveries of high Chern number QAH effect in pentalayer and tetralayer rhombohedral graphene highlights the intriguing interplay between strong interactions and spin-orbit coupling (SOC). Here we propose a minimal interacting model for spin-orbit coupled rhombohedral graphene and use the Hartree-Fock analysis to explore the phase diagram at charge neutrality. We find that with Ising SOC on one outmost graphene layer, the in-plane layer-antiferromagnetic order is the insulating ground state without displacement field. Upon increasing the gate displacement field, we find that the QAH state with Chern number being equal to the layer number emerges between layer-antiferromagentic state and layer-polarized state, which is consistent with experimental observations. We further study phase diagram for different thickness and find pentalayer is optimal for the QAH effect. Finally, we predict that QAH state is enlarged by engineering opposite Ising SOC on the opposite outmost layers of rhombohedral graphene.
@@
The quantum anomalous Hall (QAH) effect, a prominent outcome of band topology, has attracted significant attentions in condensed matter physics Chang et al. (2023); Tokura et al. (2019); Wang and Zhang (2017). The manifestation of QAH effect necessitates the breaking of time-reversal symmetry. For QAH effect discovered in magnetic topological insulators, such symmetry breaking originates from the long-range magnetic ordering of local moments Chang et al. (2013); Checkelsky et al. (2014); Kou et al. (2014); Bestwick et al. (2015); Mogi et al. (2015); Deng et al. (2020). While in the renowned Haldane model, the time-reversal breaking arises from a complex hopping term induced by the staggered magnetic flux Haldane (1988). In practical scenarios, such a complex hopping can be generated by electron interaction or spin-orbital coupling (SOC) Kane and Mele (2005); Raghu et al. (2008). However, in terms of monolayer graphene, both the interaction and SOC are too weak to actualize QAH effect. The moiré system could overcome this limitation, for electron interaction dominates over the quenched kinetic energy, exemplified in the discovery of integer and fractional Chern insulators Spanton et al. (2018); Serlin et al. (2020); Chen et al. (2020); Xie et al. (2021); Li et al. (2021); Zhao et al. (2024); Cai et al. (2023); Zeng et al. (2023); Park et al. (2023); Xu et al. (2023); Lu et al. (2023).
@@
Meanwhile, correlated insulators and QAH effect have been observed in ABC-stacked rhombohedral multilayer graphene (RMG) without any moiré patten Myhro et al. (2018); Shi et al. (2020); Kerelsky et al. (2021); Han et al. (2023a); Liu et al. (2023); Zhou et al. (2023a); Han et al. (2023b); Sha et al. (2023); Zhou et al. (2023b). Near charge neutrality, the low-energy electrons of RMG are characterized by an isospin index that includes valley and spin Grüneis et al. (2008); Koshino and McCann (2009); Zhang et al. (2010); Jung and MacDonald (2013). The bulk orbitals are dimerized, akin to the Su-Schrieffer-Heeger chain. The wave functions of the two bands closest to the Fermi level reside mostly on the non-dimerized sites in the outmost layers, with a $k^{N}$ energy dispersion for $N$ layers at each valley. These flat bands have large density of states and carry a large valley-dependent Berry phase, which results in correlated and topological states. For instance, superconductivity is observed in doped RMG Zhou et al. (2021a, b, 2022); Pantaleón et al. (2023); Zhang et al. (2023). At charge neutrality, an energy gap is spontaneously generated by Coulomb interaction Min et al. (2008); Vafek and Yang (2010); Lemonik et al. (2010); Nandkishore and Levitov (2010); Zhang et al. (2011); Jung et al. (2011); Scherer et al. (2012); Pamuk et al. (2017). Theoretically, these competing symmetry-breaking gapped states exhibit quantized anomalous, spin or valley Hall effect, which can be classified according to their Hall conductivity Nandkishore and Levitov (2010); Zhang et al. (2011); Jung et al. (2011). Experimentally, an insulating ground state without any flavor Hall effect was found at charge neutrality in rhombohedral tetralayer and pentalyer graphene Kerelsky et al. (2021); Han et al. (2023a); Liu et al. (2023); Zhou et al. (2023a). This correlated state is expected to have layer antiferromagnetic (LAF) order, where its band gap continuously decreases and eventually transitions to layer-polarized (LP) state as the gate displacement field increases. Intriguingly, the QAH state emerges between LAF and LP state in the presence of Ising SOC from nearby transition metal dichalcogenide (TMD) Gmitra and Fabian (2015, 2017); Li and Koshino (2019); Naimer et al. (2021); Zollner et al. (2023), where Chern number is $\mathcal{C}=4$ and $5$ for tetralayer and pentalyer Han et al. (2023b); Sha et al. (2023), respectively. These remarkable observations naturally lead to important questions. What is the minimal microscopic interacting model for describing the phase diagram? How is the layer dependence of the QAH effect in RMG? Why zero field QAH effect was observed only at negative displacement field in pentalayer Han et al. (2023b)? How to enlarge QAH state in the phase space?
@@
In this Letter, inspired by these experimental observations Han et al. (2023b); Sha et al. (2023), we propose a minimal interacting model for RMG with proximity Ising SOC, and investigate the phase diagram at charge neutrality using self-consistent Hartree-Fock calculations. The interactions include the short-range repulsive density-density interaction and ferromagnetic inter-valley Hund’s coupling You and Vishwanath (2022); Lu et al. (2022); Zhumagulov et al. (2023a, b). Various symmetry-broken states are degenerated at zero displacement field with the density interaction only Nandkishore and Levitov (2010); Jung et al. (2011); sup , therefore Hund’s coupling is essential to break this degeneracy. The interaction strength was determined to be consistent with experimental observations, such as band gaps and qualitative features of the phase diagram. We find that with Ising SOC, LAF-$x$ is the insulating ground state without displacement field, which breaks the spin $SU(2)$ symmetry and exhibits gapless Goldstone mode. As the displacement field increases, the phase transition occurs from LAF-$x$ to semimetal with preserved $SU(2)$ symmetry, and then turns into QAH state. We further study phase diagram for different thickness and find pentalayer is optimal for the QAH effect. Finally, we predict that QAH state is enlarged by engineering opposite Ising SOC on the top/bottom layer of RMG.
@@
$\mathcal{H}_{0}$ is the single particle tight-binding Hamiltonian Grüneis et al. (2008); Koshino and McCann (2009); Zhang et al. (2010); Jung and MacDonald (2013) of the $N$-layer rhombohedral graphene
@@
where $\tau=\pm 1$ represents $K$ and $K^{\prime}$ valley respectively, $s$ is spin index, and $c_{\mathbf{k},\tau,s}=(c_{A_{1}},c_{B_{1}},\cdots,c_{B_{N}})_{\mathbf{k},\tau,s}^{T}$ is the $2N$-component spinor including all the sublattices, with $A_{j}$ and $B_{j}$ denoting the sublattices in $j$-th layer. The explicit form for $h_{\tau}(\mathbf{k})$ are presented in Supplementary Material sup , where the parameters are adopted from Refs. Zibrov et al. (2018); Zhou et al. (2021a); Ghazaryan et al. (2023). Explicitly the low energy effective $\mathbf{k}\cdot\mathbf{p}$ model of $\mathcal{H}_{0}$ is $\mathcal{H}_{N}(\mathbf{k})\propto k^{N}\left[\cos(N\varphi_{\mathbf{k}})\varrho_{x}+\sin(N\varphi_{\mathbf{k}})\varrho_{y}\right]$, where $\cos\varphi_{\mathbf{k}}=\tau k_{x}/k$ and $\sin\varphi_{\mathbf{k}}=k_{y}/k$, and $\bm{\varrho}$ are Pauli matrices acting on layer pseudospin degree of freedom ($A_{1}/B_{N}$). The gate displacement field $\mathcal{D}$ would introduce a staggered layer potential between top ($\Delta$) and bottom ($-\Delta$) layers, which would generate a finite gap, as the low energy conduction and valance bands resides mostly in the outermost layers. $\Delta\equiv e\mathcal{D}d/(2\epsilon_{r})$, where $e$ is the electric charge, $d$ is the thickness of RMG, and $\epsilon_{r}\approx 4.5$ is the dielectric constant of hBN substrate.
@@
The nearby TMD breaks inversion symmetry and introduce Ising SOC in the bottom graphene layer through proximity effect, which breaks spin $SU(2)$ symmetry to $U(1)$. The previous ab initio calculations show that only the neighboring graphene layer could feel the proximitized Ising SOC, thus
$\mathcal{H}_{\text{soc}}=\lambda\sum_{\mathbf{k},\tau,s=\pm 1}\tau sc^{\dagger}_{\mathbf{k},\tau,s,N}c_{\mathbf{k},\tau,s,N},$ (3)
@@
where $c_{\mathbf{k},\tau,s,N}$ is electron operator on the $N$-th layer. In terms of the low energy effective model, the SOC term can be written as $\lambda\tau_{z}s_{z}(\varrho_{0}-\varrho_{z})/2$. The Rashba SOC couples two sublattices on the same layer and is negligible for the low energy physics, since a $N$-step process requires electrons hopping between low-energy sites in $A_{1}$ and $B_{N}$ via high-energy states.
The interacting Hamiltonian is modeled by the short range repulsive interactions You and Vishwanath (2022); Lu et al. (2022); Zhumagulov et al. (2023a, b),
$\mathcal{H}_{\text{int}}=\int d^{2}x\sum_{i}\frac{U}{2}n_{i}^{2}+Vn_{i,K}n_{i,K^{\prime}}-J_{\text{H}}\mathbf{S}_{i,K}\cdot\mathbf{S}_{i,K^{\prime}}$ (4)
@@
where the summation runs over all sublattices $i$, $n_{i}\equiv\sum_{\tau,s}n_{i,\tau,s}$, and $n_{i,\tau,s}$ is the electron density operator for $i$-th sublattice, $s$ spin and $\tau$ valley. The local spin density operator is $\mathbf{S}_{i,\tau}=\sum_{s,s^{\prime}}c^{\dagger}_{i,\tau,s}\mathbf{s}c_{i,\tau,s^{\prime}}$. The first term is the spin-valley $SU(4)$ invariant repulsive interaction, and $V$ represents the difference between the intra-valley and inter-valley repulsions which breaks $SU(4)$ symmetry into $SO(4)\simeq SU(2)^{K}_{s}\times SU(2)^{K^{\prime}}_{s}$, and the last term is the Hund’s coupling which further reduce the symmetry into total spin $SU(2)_{s}$ symmetry. We restrict our model to local onsite terms due to large separation between the low-energy sites $A_{1}$ and $B_{N}$. As a result, it is justified to exclude the interlayer and inter-sublattice interactions. The Hamiltonian was solved by the self-consistent Hartree-Fock methods. For numerical calculations, we choose $U=40~{}\text{eV}\cdot\mathcal{A}_{u}$, $V=-U/5$ and the momentum cut-off $\Lambda=0.16/a_{0}$, with $a_{0}$ the lattice constant of graphene and $\mathcal{A}_{u}$ as area of the unit cell. More numerical results for other parameters are listed in Supplementary Material sup .
@@
Analysis. We start by analyze all possible broken symmetry states before presenting the numerical results. By focusing on the low-energy physics, these broken symmetry states are characterized by the mass term (i.e., the potential difference between $A_{1}$ and $B_{N}$) chosen separately for ($K\uparrow$), ($K\downarrow$), ($K^{\prime}\uparrow$), and ($K^{\prime}\downarrow$) flavors Nandkishore and Levitov (2010); Zhang et al. (2011); Jung et al. (2011). Without breaking spin $SU(2)_{s}$ and valley $U(1)_{v}$ symmetry, there are $16$ distinct states depending on the sign of mass term, $m\varrho_{z}$. We categorize these 16 states into three distinct groups according to their net charge polarization, which is proportional to the sum over the spin valley of the sign of $m$. The first group exhibits full layer polarization (LP), characterized by the four flavors has the same sign of mass term. The second group has partial layer polarization (QAH with $\mathcal{C}=N$) with the sign of one flavor is opposite that of the others. And the third group has zero polarization and are likely to be the ground state in the absence of gate displacement field, which includes QAH state ($\mathcal{C}=2N$) with mass term $\tau_{z}\varrho_{z}$, LAF state with mass term $s_{z}\varrho_{z}$ and quantum spin Hall (QSH) state with mass $\tau_{z}s_{z}\varrho_{z}$.
@@
For the interacting model in Eq. (4), when $J_{\text{H}}=0$, all the three states without charge polarization are degenerate. While the presence of Ising SOC, which contains the QSH mass, may trigger the imbalance into the QSH ground state with helical edge states. However, experimentally, only the trivial correlated insulating state was found with and without SOC when external potential $\Delta$ equals to zero Sha et al. (2023); Han et al. (2023a, b); Liu et al. (2023), which indicates LAF ground state and a finite ferromagnetic Hund’s coupling ($J_{\text{H}}>0$). With Ising SOC, two specific LAF states are possible, namely, the out-of-plane ordered LAF-$z$ and in-plane ordered LAF-$x$. As we noticed that the Ising SOC term, denoted as $\lambda\tau_{z}s_{z}(\rho_{0}-\rho_{z})/2$, not only introduces a gap opening mass term, but also includes a flavor dependent chemical potential. Therefore, as illustrated in Fig. 1(b), a non-trivial band edge alignment was realized. For example, the conduction band edge of $(K,\uparrow)$ valley is aligned with the valance band edge of $(K,\downarrow)$ and $(K^{\prime},\uparrow)$ valleys, where both of them are polarized to the top layer. Then two states are possible ground state, one is planar LAF-$x$ with mass $\tau_{z}s_{x/y}\rho_{z}$, and the other is inter-valley coherence state with mass $\tau_{x/y}s_{0,z}\rho_{z}$. Numerically, we found that LAF-$x$ is favorable.
@@
Phase diagram. We now explore the phase diagram by tuning $\lambda$, $\Delta$, and $J_{\text{H}}$. Experimentally, the proximate Ising SOC strength, on the order of $1$ meV, can be varied by changing the TMD materials and also by adjusting the twisting angle between graphene and TMDLi and Koshino (2019); Zollner et al. (2023). Fig. 1(c) is the Hartree-Fock phase diagram of pentalayer graphene, showing the band gap $E_{g}$ with respect to $\lambda$ and $\Delta$, for $J_{\text{H}}=10~{}\text{eV}\cdot\mathcal{A}_{u}$. For $\Delta$ smaller than the critical value, the system exhibits LAF-$x$ order. As $\Delta$ increases, a weak first-order phase transition from LAF-$x$ state to LAF-$z$ state occurs. The phase transition happens when LAF-$z$ is semi-metallic (SM), while LAF-$x$ still maintains a tinny gap. Meanwhile, the gap of LAF-$x$ state is independent of the SOC strength, which is in agreement with the above band structure analysis. The critical potential $\Delta$ for LAF to SM transition, corresponding to the gate displacement field with $\mathcal{D}_{c}\approx 0.1$ V/nm, is consistent with the experimental observations Han et al. (2023a). By further increasing $\Delta$, Ising SOC term acts as a Haldane mass and $\Delta$ serves as the sublattice potential difference in the monolayer graphene. Then, the QAH state emerges between the LAF state and LP state. This precisely realized the Haldane model for QAH Haldane (1988). And LAF-$x$ state resembles the planar-AFM state in the Kane-Mele-Hubbard model Hohenadler et al. (2012).
@@
For experimental relevance, we limit the numerical calculation of the phase diagram with $\lambda$ value up to $2$ meV. However, for significantly larger, albeit spurious, $\lambda$ value, one could argue, based on the band edge alignment shown in Fig. 1(b), that LAF-$x$ state would undergo a crossover to the surface in-plane ferromagnetic state. Moreover, due to lack of inversion symmetry of Ising SOC, the phase diagram exhibits asymmetry with respect to $\Delta$. Specifically, the QAH state occupies a significant part in the phase diagram when $\Delta$ is less than zero, compared to its small occupation when $\Delta$ is greater than zero. Additionally, the critical value of $\lambda$ is lower when $\Delta<0$ than that when $\Delta>0$. This is consistent with the experimental observations, as the QAH effect has only been observed on one side of the displacement field for pentalayer graphene Han et al. (2023b).
@@
Let’s further study the impact of Hund’s coupling. The phase diagram in ($\Delta,J_{\text{H}}$) plane is depicted in Fig. 2(a) for $\lambda=1$ meV. The phase boundary between LAF-$x$ state and SM state disperses linearly in $J_{\text{H}}$. To understand this, we noticed the mean field decomposition of Eq. (4), for LAF state with either ordering directions, contains the mass term as $-J_{\text{H}}\mathbf{S}_{i,K}\cdot\mathbf{S}_{i,K^{\prime}}\rightarrow-J_{\text{H}}(\mathbf{S}_{i,K}\cdot\langle\mathbf{S}_{i,K^{\prime}}\rangle+\langle\mathbf{S}_{i,K}\rangle\cdot\mathbf{S}_{i,K^{\prime}})$. Therefore, for $\Delta=0$, the gap of LAF state is expected to increase approximately linearly with $J_{\text{H}}$. With the finite potential energy $\Delta$, we can naturally obtain a linear phase boundary.

### v2

@@ Abstract
The quantum anomalous Hall (QAH) effect, first proposed in the Haldane model, is a paradigmatic example of the application of band topology in condensed matter physics. The recent experimental discoveries of high Chern number QAH effect in pentalayer and tetralayer rhombohedral graphene highlight the intriguing interplay between strong interactions and spin-orbit coupling (SOC). Here we propose a minimal interacting model for spin-orbit-coupled rhombohedral graphene and use the Hartree-Fock analysis to explore the phase diagram at charge neutrality. We find that with Ising SOC on one outmost graphene layer, the in-plane layer-antiferromagnetic order is the insulating ground state without displacement field. Upon increasing the gate displacement field, we find that the QAH state with Chern number being equal to the layer number emerges between layer-antiferromagnetic state and layer-polarized state, which is consistent with experimental observations. Remarkably, we study the phase diagram for different thicknesses and find pentalayer is optimal for the QAH effect. Finally, we propose that the QAH state is enlarged by engineering opposite Ising SOC on the opposite outmost layers of rhombohedral graphene. These results will facilitate the realization of QAH states in rhombohedral graphene with different thicknesses. Our work serves as a foundation for further exploration of correlated physics of insulating state in rhombohedral graphene.
@@
The quantum anomalous Hall (QAH) effect, a prominent outcome of band topology, has attracted significant attention in condensed matter physics [1, 2, 3]. The manifestation of the QAH effect necessitates the breaking of time-reversal symmetry. For the QAH effect discovered in magnetic topological insulators, such symmetry breaking originates from the long-range magnetic ordering of local moments [4, 5, 6, 7, 8, 9]. In the renowned Haldane model, the time-reversal breaking arises from a complex hopping term induced by the staggered magnetic flux [10]. In practical scenarios, such a complex hopping can be generated by electron interaction or spin-orbit coupling (SOC) [11, 12]. However, for monolayer graphene, both the interaction and SOC are too weak to actualize the QAH effect. The moiré system could overcome this limitation, for electron interactions dominate over the quenched kinetic energy, exemplified by the discovery of integer and fractional Chern insulators  [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Moreover, the high Chern number QAH states are exceptionally rare in nature. Their realization could pave the way for the development of next-generation, low-dissipation electronic devices [25, 26].
@@
Meanwhile, correlated insulators and the QAH effect have been observed in ABC-stacked rhombohedral multilayer graphene (RMG) without any moiré patten [27, 28, 29, 30, 31, 32, 33, 34, 35]. Near charge neutrality, the low-energy electrons of RMG are characterized by an isospin index that includes valley and spin [36, 37, 38, 39]. The bulk orbitals are dimerized, akin to the Su-Schrieffer-Heeger chain. The wavefunctions of the two bands closest to the Fermi level reside mostly on the non-dimerized sites in the outmost layers, with a $k^{N}$ energy dispersion for $N$ layers at each valley. These flat bands have large density of states and carry a large valley-dependent Berry phase, which results in correlated and topological states. For instance, superconductivity is observed in doped RMG [40, 41, 42, 43, 44]. At charge neutrality, an energy gap is spontaneously generated by Coulomb interaction  [45, 46, 47, 48, 49, 50, 51, 52]. Theoretically, these competing symmetry-breaking gapped states exhibit quantized anomalous, spin or valley Hall effects, which can be classified according to their Hall conductivity  [48, 49, 50]. Experimentally, an insulating ground state without any flavor Hall effect was found at charge neutrality in rhombohedral tetralayer and pentalyer graphene [29, 30, 31, 32]. This correlated state is expected to have layer antiferromagnetic (LAF) order, where its band gap continuously decreases and eventually transitions to the layer-polarized (LP) state as the gate displacement field increases. Intriguingly, the QAH state emerges between the LAF and LP state in the presence of Ising SOC from nearby transition metal dichalcogenide (TMD) [53, 54, 55, 56, 57], where Chern number is $\mathcal{C}=4$ and $5$ for tetralayer and pentalyer [33, 34], respectively. These remarkable observations naturally lead to important questions. What is the minimal microscopic interacting model for describing the phase diagram? How is the layer dependence of the QAH effect in RMG? Why zero-field QAH effect was observed only at negative displacement field in pentalayer [33]? How to enlarge the QAH state in the phase space?
@@
In this Letter, inspired by these experimental observations [33, 34], we answer these questions by proposing a minimal interacting model for proximitized RMG, and investigating the phase diagram at charge neutrality using self-consistent Hartree-Fock calculations. The interactions include the short-range repulsive density-density interaction and ferromagnetic inter-valley Hund’s coupling [58, 59, 60, 61, 62, 63, 64]. The interaction strength was determined to be consistent with the experimental results. We study the phase diagram for different thicknesses and find the pentalayer is optimal for the QAH effect, which explains well the experiments. The good agreement between our calculation and experimental results strongly validates our model. We further propose that the QAH state is enlarged by engineering opposite Ising SOC on the top and bottom layers of RMG.
@@
$\mathcal{H}_{0}$ is the single particle tight-binding Hamiltonian [36, 37, 38, 39] of the $N$-layer rhombohedral graphene
@@
where $\tau=\pm 1$ represents $K$ and $K^{\prime}$ valley respectively, $s$ is spin index, and $c_{\mathbf{k},\tau,s}=(c_{A_{1},\mathbf{k},\tau,s},c_{B_{1},\mathbf{k},\tau,s},\cdots,c_{B_{N},\mathbf{k},\tau,s})^{T}$ is the $2N$-component spinor including all the sublattices, with $A_{j}$ and $B_{j}$ denoting the sublattices in $j$-th layer. Taking the tetralayer system as an example, the Hamiltonian is
$h_{\tau,\text{4L}}=\begin{pmatrix}h_{0}(1)&h_{1}&h_{2}&0\\ h^{\dagger}_{1}&h_{0}(2)&h_{1}&h_{2}\\ h^{\dagger}_{2}&h^{\dagger}_{1}&h_{0}(3)&h_{1}\\ 0&h^{\dagger}_{2}&h^{\dagger}_{1}&h_{0}(4)\\ \end{pmatrix},$ (3)
where
$h_{\tau,0}(i)$ $=$ $\begin{pmatrix}\delta(2i-1)&v_{0}\pi^{\dagger}\\ v_{0}\pi&\delta(2i)\\ \end{pmatrix},\quad h_{\tau,1}=\begin{pmatrix}v_{4}\pi^{\dagger}&v_{3}\pi\\ \gamma_{1}&v_{4}\pi^{\dagger}\\ \end{pmatrix},$
$h_{\tau,2}$ $=$ $\begin{pmatrix}0&\gamma_{2}/2\\ 0&0\\ \end{pmatrix}.$ (4)
@@
and $\pi=\tau k_{x}+ik_{y}$, $\tau=\pm 1$ for $K$ and $K^{\prime}$ valley. The $\delta(i)=\delta$ for $i=1$ or $2N$ and equals to zero for others. And $v_{i}=\sqrt{3}a\gamma_{i}/2$, with $a=0.246$ nm as the lattice constant of graphene. The parameters are chosen from Refs. [40, 65, 66] with $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\delta)=(3.1,0.38,-0.015,-0.29,-0.141,0.0105)~{}\text{eV}$. Explicitly the low energy effective $\mathbf{k}\cdot\mathbf{p}$ model of $\mathcal{H}_{0}$ is $\mathcal{H}_{N}(\mathbf{k})\propto k^{N}\left[\cos(N\varphi_{\mathbf{k}})\varrho_{x}+\sin(N\varphi_{\mathbf{k}})\varrho_{y}\right]$, where $\cos\varphi_{\mathbf{k}}=\tau k_{x}/k$ and $\sin\varphi_{\mathbf{k}}=k_{y}/k$, and $\bm{\varrho}$ are Pauli matrices acting on layer pseudospin degree of freedom ($A_{1}/B_{N}$). The gate displacement field $\mathcal{D}$ would introduce a interlayer potential between top ($\Delta$) and bottom ($-\Delta$) layers, which would generate a finite gap, as the low energy conduction and valance bands resides mostly in the outmost layers. $\Delta\equiv e\mathcal{D}d/(2\epsilon_{r})$, where $e$ is the electric charge, $d$ is the thickness of RMG, and $\epsilon_{r}\approx 4.5$ is the dielectric constant of hBN substrate [63].
@@
The nearby TMD breaks inversion symmetry and introduces Ising SOC in the bottom graphene layer through proximity effect, which breaks the spin $SU(2)$ symmetry to $U(1)$. The previous ab initio calculations show that only the neighboring graphene layer could feel the proximitized Ising SOC, thus
$\mathcal{H}_{\text{soc}}=\lambda\sum_{\mathbf{k},\tau,s=\pm 1}\tau sc^{\dagger}_{\eta_{N},\mathbf{k},\tau,s}c_{\eta_{N},\mathbf{k},\tau,s},$ (5)
@@
where $c_{\eta_{N},\mathbf{k},\tau,s}$ is electron operator on the $N$-th layer with $\eta=A/B$ for sublattices. In terms of the low energy effective model, the SOC term can be written as $\lambda\tau_{z}s_{z}(\varrho_{0}-\varrho_{z})/2$. The Rashba SOC couples two sublattices on the same layer and is negligible for the low energy physics, since a $N$-step process requires electrons hopping between low-energy sites in $A_{1}$ and $B_{N}$ via high-energy states.
The interacting Hamiltonian is modeled by the short range repulsive interactions [58, 59, 60, 61, 62],
$\mathcal{H}_{\text{int}}=\int d^{2}x\sum_{i}(\frac{U}{2}n_{i}^{2}+Vn_{i,K}n_{i,K^{\prime}}-J_{\text{H}}\mathbf{S}_{i,K}\cdot\mathbf{S}_{i,K^{\prime}})$ (6)
@@
where the summation runs over all sublattices $i$, $n_{i}\equiv\sum_{\tau,s}n_{i,\tau,s}$, and $n_{i,\tau,s}$ is the electron density operator for $i$-th sublattice, $s$ spin and $\tau$ valley. The local spin density operator is $\mathbf{S}_{i,\tau}=\sum_{s,s^{\prime}}c^{\dagger}_{i,\tau,s}\mathbf{s}c_{i,\tau,s^{\prime}}$. The first term is the spin-valley $SU(4)$ invariant repulsive interaction, and $V$ represents the difference between the intra-valley and inter-valley repulsions which breaks $SU(4)$ symmetry into $SO(4)\simeq SU(2)^{K}_{s}\times SU(2)^{K^{\prime}}_{s}$, and the last term is the Hund’s coupling which further reduce the symmetry into total spin $SU(2)_{s}$ symmetry. We restrict our model to local onsite terms due to large separation between the low-energy sites $A_{1}$ and $B_{N}$. As a result, it is justified to exclude the interlayer and inter-sublattice interactions. The Hamiltonian was solved by the self-consistent Hartree-Fock methods. For numerical calculations, we choose $U=40~{}\text{eV}\cdot\mathcal{A}_{u}$, $V=-U/5$ and the momentum cut-off $\Lambda=0.16/a_{0}$, with $a_{0}$ the lattice constant of graphene and $\mathcal{A}_{u}$ as area of the unit cell. More numerical results for other parameters are listed in Supplementary Material [64].
@@
Analysis. We start by analyzing all possible broken symmetry states before presenting the numerical results. By focusing on the low-energy physics, these broken symmetry states are characterized by the mass term (i.e., the potential difference between $A_{1}$ and $B_{N}$) chosen separately for ($K\uparrow$), ($K\downarrow$), ($K^{\prime}\uparrow$), and ($K^{\prime}\downarrow$) flavors [48, 49, 50]. Without breaking spin $SU(2)_{s}$ and valley $U(1)_{v}$ symmetries, there are $16$ distinct states depending on the sign of the mass term, $m\varrho_{z}$. We categorize these 16 states into three distinct groups according to their net layer polarization, which is proportional to the sum over the spin valley of the sign of $m$. The first group exhibits full layer polarization (LP), characterized by the four flavors has the same sign of mass term. The second group has partial layer polarization (QAH with $\mathcal{C}=N$) with the sign of one flavor is opposite that of the others. And the third group has zero polarization and is likely to be the ground state in the absence of a gate displacement field, which includes the QAH state ($\mathcal{C}=2N$) with mass term $\tau_{z}\varrho_{z}$, the LAF state with mass term $s_{z}\varrho_{z}$ and the quantum spin Hall (QSH) state with mass $\tau_{z}s_{z}\varrho_{z}$.
@@
For the interacting model in Eq. (6), when $V=0$ and $J_{\text{H}}=0$, all the three states without layer polarization are degenerate. While the presence of Ising SOC, which contains the QSH mass, may trigger the imbalance into the QSH ground state with helical edge states. However, experimentally, only the trivial correlated insulating state was found with and without SOC when the external potential $\Delta$ equals zero [34, 30, 33, 31], which indicates an LAF ground state and a finite ferromagnetic Hund’s coupling ($J_{\text{H}}>0$). With Ising SOC, two specific LAF states are possible, namely, the out-of-plane ordered LAF-$z$ and in-plane ordered LAF-$x$. As we noticed that the Ising SOC term, denoted as $\lambda\tau_{z}s_{z}(\varrho_{0}-\varrho_{z})/2$, not only introduces a gap opening mass term, but also includes a flavor dependent chemical potential. Therefore, as illustrated in Fig. 1(b), a non-trivial band edge alignment was realized. For example, the valence band edge of $(K,\uparrow)$ valley is aligned with the conduction band edge of $(K,\downarrow)$ and $(K^{\prime},\uparrow)$ valleys, where both of them are polarized to the top layer. Then two states are possible ground state, one is planar LAF-$x$ with mass $\tau_{z}s_{x/y}\varrho_{z}$, and the other is inter-valley coherence state with mass $\tau_{x/y}s_{z}\varrho_{z}$. Numerically, we found that LAF-$x$ is favorable.
@@
Phase diagram. We now explore the phase diagram by tuning $\lambda$, $\Delta$, and $J_{\text{H}}$. Experimentally, the proximate Ising SOC strength, on the order of $1$ meV, can be varied by changing the TMD materials and also by adjusting the twisting angle between graphene and TMD[55, 56, 57]. Fig. 1(c) is the Hartree-Fock phase diagram of pentalayer graphene, showing the band gap $E_{g}$ with respect to $\lambda$ and $\Delta$, for $J_{\text{H}}=10~{}\text{eV}\cdot\mathcal{A}_{u}$. For $\Delta$ smaller than the critical value of the LAF state to semi-metal transition (about $15$ meV), the system exhibits LAF-$x$ order. As $\Delta$ increases, a weak first-order phase transition from the LAF-$x$ state to the LAF-$z$ state occurs where the band gap changes about $0.5$ meV [64]. The phase transition happens when LAF-$z$ is semi-metallic (SM), while LAF-$x$ still maintains a tiny gap. The SM state is a compensated semi-metal with electron and hole pockets. Meanwhile, the gap of LAF-$x$ state is independent of the SOC strength, which is in agreement with the above band structure analysis. The critical potential $\Delta$ for the LAF to SM transition, corresponding to the gate displacement field with $\mathcal{D}_{c}\approx 0.1$ V/nm, is consistent with the experimental observations [30]. By further increasing $\Delta$, Ising SOC term acts as a Haldane mass and $\Delta$ serves as the sublattice potential difference in the monolayer graphene. More precisely, in the SM phase, the low energy band consists of two valleys with the same spin. Therefore, the system is effectively spinless, the same as the Haldane model. If we consider the top and bottom layers as the A/B sublattices in the honeycomb lattice, the potential $\Delta$ represents the sublattice potential difference. Since the Ising-SOC has opposite signs for the two valleys, this term indeed maps to the Haldane mass[33]. Then, the QAH state emerges between the LAF state and LP state. This precisely realized the Haldane model for QAH [10]. And the LAF-$x$ state resembles the planar-AFM state in the Kane-Mele-Hubbard model [67].
@@
For experimental relevance, we limit the numerical calculation of the phase diagram with $\lambda$ value up to $2$ meV. However, for significantly larger, albeit spurious, $\lambda$ values, one could argue, based on the band edge alignment shown in Fig. 1(b), that the LAF-$x$ state would undergo a crossover to the surface in-plane ferromagnetic state. Moreover, due to lack of inversion symmetry of Ising SOC, the phase diagram exhibits asymmetry with respect to $\Delta$. Specifically, the QAH state occupies a significant part in the phase diagram when $\Delta$ is less than zero, compared to its small occupation when $\Delta$ is greater than zero. Additionally, the critical value of $\lambda$ is lower when $\Delta<0$ than that when $\Delta>0$. This is consistent with the experimental observations, as the QAH effect has only been observed on one side of the displacement field for pentalayer graphene [33].
@@
We further study the impact of Hund’s coupling. The phase diagram in ($\Delta,J_{\text{H}}$) plane is depicted in Fig. 2(a) for $\lambda=1$ meV. The phase boundary between LAF-$x$ state and SM state disperses linearly in $J_{\text{H}}$. To understand this, we noticed that the mean field decomposition of Eq. (6), for LAF state with either ordering directions, contains the mass term as $-J_{\text{H}}\mathbf{S}_{i,K}\cdot\mathbf{S}_{i,K^{\prime}}\rightarrow-J_{\text{H}}(\mathbf{S}_{i,K}\cdot\langle\mathbf{S}_{i,K^{\prime}}\rangle+\langle\mathbf{S}_{i,K}\rangle\cdot\mathbf{S}_{i,K^{\prime}})$. Therefore, for $\Delta=0$, the gap of the LAF state is expected to increase approximately linearly with $J_{\text{H}}$. With the finite potential energy $\Delta$, we can naturally obtain a linear phase boundary.